OpenStudy (anonymous):

Let f(x) = -7 x^{4} sqrt{x} + ({2}/ {x^{3} sqrt{x} }). find f'( x )

OpenStudy (anonymous):

\[ f(x) = -7 x^{4}\sqrt{x} + (\frac{2}{ x^{3} \sqrt{x}})\] like that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

rewrite as \[-7x^{\frac{9}{2}}+2x^{-\frac{7}{2}}\] and use the power rule twice

OpenStudy (anonymous):

 Like this @satellite73